OpenStudy (anonymous):

8. The diameter of a Frisbee is 12 in. What is the area of the Frisbee? A. 18.84 sq. in. B. 113.04 sq. in. C. 452.16 sq. in. D. 37.68 sq. in.

OpenStudy (anonymous):

a=r*pi*2

OpenStudy (insa):

frisbee is a circle right? u can find the area using pi*r^2

OpenStudy (anonymous):

65.94?

OpenStudy (anonymous):

i mean pi *r^2

OpenStudy (anonymous):

what does that =@Tksmith514

OpenStudy (insa):

@DaBest21 won't it be like 3.14*6^2?

OpenStudy (anonymous):

9^2=81 so 36*3.14=what @Tksmith514

OpenStudy (anonymous):

yes srry again

OpenStudy (anonymous):

@Tksmith514 are u there

OpenStudy (anonymous):

9^2=81 so 36*3.14=what @Tksmith514

OpenStudy (anonymous):

6^2=36

OpenStudy (anonymous):

hello